# Co-Reinforcement Learning for Unified Multimodal Understanding and Generation

### v1

@@ Abstract
This paper presents a pioneering exploration of reinforcement learning (RL) via group relative policy optimization for unified multimodal large language models (ULMs), aimed at simultaneously reinforcing generation and understanding capabilities. Through systematic pilot studies, we uncover the significant potential of ULMs to enable the synergistic co-evolution of dual capabilities within a shared policy optimization framework. Building on this insight, we introduce CoRL, a co-reinforcement learning framework comprising a unified RL stage for joint optimization and a refined RL stage for task-specific enhancement. With the proposed CoRL, our resulting model, ULM-R1, achieves average improvements of 7% on three text-to-image generation datasets and 23% on nine multimodal understanding benchmarks. These results demonstrate the effectiveness of CoRL and highlight the substantial benefit of reinforcement learning in facilitating cross-task synergy and optimization for ULMs.
@@ 1 Introduction
In light of our preliminary findings, we propose CoRL, a co-evolutionary reinforcement learning framework designed to synergistically improve the understanding and generation capabilities of ULMs. CoRL follows a two-stage RL procedure: a unified RL stage for joint optimization of dual capabilities and a refined RL stage for task-specific enhancement. In the first stage, the policy ULM is optimized through a unified GRPO algorithm with diverse rewards on a carefully curated dataset spanning both understanding and generation tasks. To effectively guide policy optimization in visual generation, we introduce a bidirectional cycle consistency reward and a text-image matching reward, which together promote semantic consistency and faithfulness of synthesized images to their corresponding prompts. The designed rewards complement typical multimodal understanding rewards (i.e., accuracy and format) within a unified group, enabling cross-task joint optimization. In the subsequent stage, we independently reinforce the policy’s understanding and generation capabilities using respective rewards and tailored datasets for task-specific refinement.
@@
  • We introduce a co-evolutionary reinforcement learning framework, CoRL, to synergistically enhance the dual capabilities of ULMs using a unified-then-refined RL paradigm.
@@ 3.2 Pilot Exploration
Figure 2: Overview of CoRL, a co-evolutionary reinforcement learning framework to jointly improve the dual capabilities of ULMs. CoRL adopts a two-stage RL procedure, comprising a unified RL stage for joint optimization and a refined RL stage for task-specific enhancement.
@@ 5 Conclusion
In this work, we investigate how to jointly enhance the understanding and generation capabilities of ULMs, and propose a co-evolutionary RL framework (CoRL). Within the proposed CoRL, the policy model undergoes unified RL for joint optimization and refined RL for task-specific enhancement, yielding ULM-R1. Extensive evaluations across diverse understanding and generation benchmarks demonstrate the effectiveness of CoRL and the advantage of ULM-R1.

### v2

@@ Abstract
This paper presents a pioneering exploration of reinforcement learning (RL) via group relative policy optimization for unified multimodal large language models (ULMs), aimed at simultaneously reinforcing generation and understanding capabilities. Through systematic pilot studies, we uncover the significant potential of ULMs to enable the synergistic co-evolution of dual capabilities within a shared policy optimization framework. Building on this insight, we introduce CoRL, a co-reinforcement learning framework comprising a unified RL stage for joint optimization and a refined RL stage for task-specific enhancement. With the proposed CoRL, our resulting model, ULM-R1, achieves average improvements of 7% on three text-to-image generation datasets and 23% on nine multimodal understanding benchmarks. These results demonstrate the effectiveness of CoRL and highlight the substantial benefits of reinforcement learning in facilitating cross-task synergy and optimization for ULMs. Code is available at https://github.com/mm-vl/ULM-R1.
@@ 1 Introduction
In light of our preliminary findings, we propose CoRL, a co-reinforcement learning framework designed to synergistically improve the understanding and generation capabilities of ULMs. CoRL follows a two-stage RL procedure: a unified RL stage for joint optimization of dual capabilities and a refined RL stage for task-specific enhancement. In the first stage, the policy ULM is optimized through a unified GRPO algorithm with diverse rewards on a carefully curated dataset spanning both understanding and generation tasks. To effectively guide policy optimization in visual generation, we introduce a bidirectional cycle consistency reward and a text-image matching reward, which together promote semantic consistency and faithfulness of synthesized images to their corresponding prompts. The designed rewards complement typical multimodal understanding rewards (i.e., accuracy and format) within a unified group, enabling cross-task joint optimization. In the subsequent stage, we independently reinforce the policy’s understanding and generation capabilities using respective rewards and tailored datasets for task-specific refinement.
@@
  • We introduce a co-reinforcement learning framework, CoRL, to synergistically enhance the dual capabilities of ULMs using a unified-then-refined RL paradigm.
@@ 3.2 Pilot Exploration
Figure 2: Overview of CoRL, a co-reinforcement learning framework to jointly improve the dual capabilities of ULMs. CoRL adopts a two-stage RL procedure, comprising a unified RL stage for joint optimization and a refined RL stage for task-specific enhancement.
@@ 5 Conclusion
In this work, we investigate how to jointly enhance the understanding and generation capabilities of ULMs, and propose a co-reinforcement learning framework (CoRL). Within the proposed CoRL, the policy model undergoes unified RL for joint optimization and refined RL for task-specific enhancement, yielding ULM-R1. Extensive evaluations across diverse understanding and generation benchmarks demonstrate the effectiveness of CoRL and the advantage of ULM-R1.